Grade 11General Physics

find the extreme maximna and minima values of f(x)=sqrt of x -4

rucha patil
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear student
 
f(x) = \sqrt(x-4)
 
Now x-4 should be greater than or equal to zero.
 
Hence minimum value when x = 4
Hence
 
f(x) min = \sqrt(4-4) = 0
 
Now
 
f(x) max = infinity